### v1

# RealCQA-V2 : Visual Premise Proving thanks: WIP: Data will be uploaded soon

Abstract
We introduce Visual Premise Proving (VPP), a novel task tailored to refine the process of chart question answering by deconstructing it into a series of logical premises. Each of these premises represents an essential step in comprehending a chart’s content and deriving logical conclusions, thereby providing a granular look at a model’s reasoning abilities. This approach represents a departure from conventional accuracy-based evaluation methods, emphasizing the model’s ability to sequentially validate each premise and ideally mimic human analytical processes. A model adept at reasoning is expected to demonstrate proficiency in both data retrieval and the structural understanding of charts, suggesting a synergy between these competencies. However, in our zero-shot study using the sophisticated MATCHA model on a scientific chart question answering dataset, an intriguing pattern emerged. The model showcased superior performance in chart reasoning (27%) over chart structure (19%) and data retrieval (14%). This performance gap suggests that models might more readily generalize reasoning capabilities across datasets, benefiting from consistent mathematical and linguistic semantics, even when challenged by changes in the visual domain that complicate structure comprehension and data retrieval. Furthermore, the efficacy of using accuracy of binary QA for evaluating chart reasoning comes into question if models can deduce correct answers without parsing chart data or structure. VPP highlights the importance of integrating reasoning with visual comprehension to enhance model performance in chart analysis, pushing for a balanced approach in evaluating visual data interpretation capabilities.
1 Introduction
Understanding and reasoning with visual data is crucial for Multi-Modal Question Answering (MMQA) systems. However, current methods do not effectively evaluate the sequential reasoning needed for complete comprehension, resulting in a gap between visual data representation and semantic interpretation. Current benchmarks in structured visual reasoning, such as those used in [1, 2], primarily rely on one-to-one accuracy metrics, which do not adequately capture the true reasoning capabilities of these models. While formal reasoning tasks like Conjecture/Theorem Proving and First-Order Logic (FOL) Verification have been extensively studied in structured domains like mathematics and natural language processing, such formalization is largely absent in vision-language MMQA settings.
To bridge this gap, we introduce Visual Premise Proving (VPP), which formalizes reasoning in Chart Question Answering (CQA) by breaking down the reasoning process into logical premises, each representing a step needed to understand a chart and draw a conclusion. This approach shifts the emphasis from mere accuracy to validating a model’s ability to replicate human-like analytical processes through step-by-step sequential premise validation.
Our contributions are threefold:
  • We introduce the Visual Premise Proving (VPP) task to formalize reasoning sequences in the domain of Chart Question Answering.
  • We propose two novel evaluation metrics, $ACC_{VPP}$ and $DCP$, to provide interpretable and explainable assessments of visual reasoning capabilities.
  • We curate a challenging dataset comprising 10 million real scientific chart question-answer pairs, annotated with premise-conclusion sequences, for evaluating sequential reasoning.
1.1 Problem Definition: Visual Reasoning Chains
The formalization of generalized visual reasoning is still in its early stages and involves tackling complex, higher-order reasoning problems. To make this problem tractable, we introduce two key constraints:
First, we limit the scope to First-Order Logic (FOL) problems. FOL verification is inherently complex and often NP-hard, particularly when dealing with natural language, which have inherent lexical, syntactic, and contextual ambiguities unless manually annotated.
Second, we focus specifically on Chart Question Answering (CQA). The recently released RealCQA dataset [3] serves as an ideal testbed for this task. This dataset, which is template-based (avoiding complex language), manually annotated (ensuring consistency and completeness), and grounded in mathematical logic (ensuring coherent quantification), makes the process of logical verification more manageable.
By restricting our problem to this structured and controlled environment, we establish a foundational framework for formalizing visual reasoning in MMQA. Our current work begins with chart QA and aims to extend to general visual question answering (VQA) and, eventually, more abstract visual reasoning in future works. The primary challenge lies in identifying and verifying visual premises and reasoning chains in diverse and unconstrained environments. Our work takes a step in this direction by focusing on the constrained domain of Chart Question Answering, where we introduce the VPP task to evaluate a model’s reasoning capacity in a manner analogous to human chart analysis. This not only measures the model’s final answer accuracy but also its ability to replicate the reasoning process across multiple steps.
2 Background
We discuss some recent works proposed for multi-modal document understanding domain. Then further specific Chart Reasoning tasks, datasets and models.
2.1 Vision-Language Models
Early models, such as LayoutLM [4], used separate encoders for vision and language, combining their outputs at a later stage (late fusion) to generate text from image and text inputs. Dessurt [5] proposed a novel token processing mechanism to provide OCR free document processing. Pix2Struct [6], similarly extend the VIT architecture with combined fixed resolution image and text tokens as input and showcase larger generalizability by converting web page screenshots to HTML. These advancements signal a shift towards more integrated and efficient multi-modal methods. Recently, the emergence of large vision-language models (LVLMs) like LLAVA, Gemini, GPT-4 etc [7, 8, 9] leverage similar novel token and early to late fusion strategies for combining visual input with large language models.
Figure 1: Diverse levels of visual reasoning are required for robust chart analysis.
2.2 Visual Reasoning over Charts
Charts convey complex information through visual elements like trend lines and legends, requiring detailed interpretation of various visual tasks (see Figure 1). These require prediction and evaluation at each step and are referred to as ‘Dense’ in Table1. Chart-to-text tasks like QA and summarizing involve directly evaluating chart understanding tasks through the text output.
Dataset # Img Source Task
FigureSeer [10] 1k ArXiv Dense
UB-PMC [11] 28k PubMed Central Dense
Real CQA [3] 28k PubMed Central QA
ChartQA [1] 22k Pew/Statista/OWID QA
C2T [12] 82k Pew/Statista Summary
EC400k [13] 400k Excel Dense
FQA [14] 180k Synthetic QA
DVQA [14] 300k Synthetic QA
LeafQA [15] 200k Synthetic QA
Table 1: Popular Chart Datasets, Sources and Tasks
Figure 2: Deconstructing Chart QA to hierarchical premises.
2.2.1 Chart Datasets
Synthetic datasets like FQA, DVQA, and LeafQA, have extensive scale (180k to 300k charts) and dense text annotations(components, captions, summaries, question answers), usually leverage real world tabular data-source and plot images using standard libraries like Matplotlib.
Real-world chart datasets such as, FigureSeer, UB-PMC, and EC400k, are costly to annotate and have limited number of images/annoations available. Scientific charts, from sources like ArXiV and PMC encompass a wide range of technical and stochastic data required for academic discourse as compared to business oriented excel charts.
While synthetic charts are easy to scale they lack fidelity with real-world chart images and under-perform with even a slight variation in the data distribution. Digital-born scientific publications, especially pose a significant challenge due to their complex visual layouts and intricate details, such as dense plot elements, noisy overlapping lines and bars, high concentration of math and special symbols etc.
FigureSeer [10], consists of $\sim$1k densely annotated line charts from arXiv publications. RealCQA consists of $\sim$ 2Mil QA pairs based on 240 templates for $\sim$28k human annotated charts first proposed by UB-PMC from pubmed central publications. EC400k provides line, bar and pie plots from business based sources and obfuscates text in charts, making any further semantic use impossible. ChartQA, C2T used for chart to text tasks are primarily based on more straight forward data from sources like Pew, Statista and Our World in Data(OWID).
We base our study on RealCQA dataset as it is the only dataset that provides a combination of challenging real-world data, scientific chart images, and dense annotations of both structural elements and textual information to ensure the creation of verifiable FOL reasoning sequences. Two requirements for creating a valid FOL are (i) a closed set of variables and (ii) a closed set of predicates. The closed set of variables includes chart components such as tick values, axis titles, legend labels, etc manually identified for the chart structure prediction task of UB-PMC. Further the QA templates used for RealCQA were handcrafted by domain experts, and generate reasoning-based questions by performing mathematical comparisons between a given subset of chart components a.k.a our variables. This ensures completeness of predicate logic. We provide exhaustive details over variables, predicates, premises, and our curated FOL sequences for each template in supplementary section.
2.2.2 Chart Understanding Models :
Dense prediction models like Cached [16] focus on chart structure prediction and proposed a custom local-global context modelling architecture based on Swin-Transformers. Chart-Ocr[13] and SpaDen [17] focuses on using keypoints for tabular data extraction, by recreating plot area components like bars, lines etc and end-to-end data extraction with legend mapping.
Works focusing on chart question answering, like CRCT [18] use classification and regression predictors based on ‘Answer Type’. ChartQA[1] introduced template and human annotated question answers, baselined through offline tabular extraction using ChartOCR and answer prediction using VLT5. Matcha [19] utilizes the Pix2Struct model[6] with training on HTML structure, Math and ChartQA to get an impressive jump in $20\%$ accuracy on the ChartQA dataset.
To tackle Chart to Text tasks, Chartllama [20] first proposed using a generic visual encoder with LLAMA. ChartReader [21] further leverages a keypoint-based chart component detector with a T5-based text decoder. StructChart [22] showed novelty by reformulating chart data into triplets, effectively capturing both the structure and semantics inherent in the data. ChartAssistant [23] integrates Swin and ViT encoders with Bart and LLAMA language decoders, evaluating across multiple datasets. All of these chart to text works suffer from the inherent ambiguity in one-to-one evaluation of QA based visual reasoning.
3 Chart FOL Premise Proving
(a) Structure Premises
(b) True/False Count
(c) PCA
(d) tSNE
Figure 3: Distribution of NLP Premises (Best viewed digital, zoom and color)
We discuss our methodology for the proposed task. First we discuss our base ideation and task formalization, then discuss some meta charecteristics of the created data from the approach and finally our proposed evaluation metric for the task.
We want to build a system for evaluation of visual reasoning from first principles as depicted in Figure 2. Our basic idea derives this theory from the way human beings read semantically arranged visual representations of data, i.e. charts. We start from the input an Image $I\in R^{3}$ and a question. Each question context is confined to the template variables and predicates. This can be systematically characterized into a series of premise-conclusions per question, visually grounded in the plot area.
3.1 VPP Task
"We start by defining a Premise-Conclusion Pair, where each premise logically leads to a conclusion. These premises are articulated in FOL, utilizing quantifiers such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$), alongside predicates applicable to these variables. Subsequently, we establish a Chain of Reasoning. This sequence involves linking each proposition to the next, ensuring that each step logically follows from the previous one as described in Algorithm 1 and the example in 3.2.1, and 3.2.2 . The task culminates in Proving the Premises, where the final premise-conclusion, the original question—answer is derived from the interconnected premises through the established reasoning chain. Successfully completing this task confirms that the conclusion is a logical consequence of the premises. Since each premise is an FOL statement we can represent our final premise conclusion as natural language strings, as well as mathematical graph representations [24] using abstract syntax trees (AST) uniquely identifying each logical sequence. Representing textual premises as graphs where nodes correspond to entities or concepts mentioned in the premises, and edges represent logical or causal relationships between these entities enables problem formulation as a graph problem for visual logic verification studied independantly of the natural language complexity. While we have released the AST representation of our dataset we currently focus on the NLP version for this work.
3.2 Premise Conclusion Sequence Creation
We define our step by step process of identifying and creating visual premises for charts in Algorithm 1. We create four types of chart premises:
  1. 1.
    Structural Premises (SP) Identify the chart’s structure, categorized into 10 types (e.g., 0: Chart Type, 1: Y-Title, 2: X-Title, 3: Y-Range, 4: X-Range, 5: Categorical, 6: X-Tick, 7: Y-Tick, 8: # of Dataseries, 9: Legend Label as shown in Fig 3(a)).
  2. 2.
    Data Premises (DP) per chart image for data retrieval verification of that plot area.
  3. 3.
    Reasoning Premises (RP) per question template requiring logical deduction based on chart data and explicit calculations like comparing or transforming data values.
  4. 4.
    Mathematical Premises (MP) for a subset of question templates consisting of implicit math; like finding Pearson’s correlation in a bi-variate scatter plot or calculating distribution tendencies like the median, upper-quartile of a box-plot etc.
Thus each question template is deconstructed to sequential premise templates /AST. Each premise template is used to generate a premise string and has true or false conclusion specific to query values of that chart used to generate the string. For True conclusions we use actual variable value present in the groundtruth chart annotation. For False conclusions we randomly use any other relevant variable values present in the current chart (eg. out of range tick value, axis title instead of legend label etc). This setup effectively breaks down the original question into a comprehensible set of binary premise-conclusions, generated in a ratio of 1:3 true-to-false for each as shown in Figure 3(b). This is done by design to bias the dataset with distractor (False) representations and prevent overfitting to actual (True) conclusion statements. In generating the final Premise strings we further add vocabulary complexity by choosing at random 1 of 3 paraphrased templates created using a pre-trained T5-Transformer from original handcrafted premise template.
The PCA and t-SNE plots of pretrained-BERT embeddings as shown in Figures 3(c), and 3(d) for 25,364 unique words indicate a highly diverse vocabulary, with a broad and evenly distributed semantic space and no apparent clustering, highlighting rich semantic coverage. In contrast to the focused and less diverse vocabularies in popular VQA tasks, which show distinct clusters around common categories and actions, this corpus encompasses a wider range of topics and semantics. The extensive diversity and rich semantic distribution suggest a more nuanced and challenging dataset.
Algorithm 1 Parsing a Chart from First Principles
Input: QA template, FOL templates
Parameters: Chart structure variables per image ($x,y,\ell,t$), QA values per string ($v,d$), FOL predicates
Output: Sequence of FOL premise-conclusion pairs
1:  while Chart components {$x,y,\ell,t,v,d$} exist do
2:     Recognize axis labels and verify variables: $\forall x\in\text{Axis Labels},\exists y\in\text{Variables}:\text{InRange}(y)$
3:     if Legend exists then
4:        Locate legend and map data series: $\exists\ell\in\text{Legends}:\text{LocateRecognize}(\ell)$
5:     end if
6:     Verify tick marks on axis: $\forall t\in\text{Tick Marks},\text{Valid}(t)\wedge\neg\text{BrokenAxis}(t)$
7:     Verify data values on plot area: $\forall v\in\text{Tick Values},\exists d\in\text{Data Points}:\text{AtLocation}(v,d)$
8:     if Legend exists then
9:        Check data values for each legend: $\forall v\in\text{Tick Values},\ell\in\text{Legends},\exists d\in\text{Data Points}:\text{AtLocationWithLegend}(v,\ell,d)$
10:     end if
11:  end while
12:  Verify predicate logic for validated data values
13:  if Implicit math transformation required then
[…]
15:  end if
3.2.1 Example Query Template
Is the difference between the value of $Y$ for legendlabel1 at the $i$-th $x$ tick and the $j$-th $x$ tick greater than the difference between the value of $Y$ for legendlabel2 at the $i$-th $x$ tick and the $j$-th $x$ tick?
3.2.2 Deconstructed Sequential Premises
[…]
3.3 VPP Evaluation
The $Acc_{VPP}$ metric evaluates a model’s ability to perform complete and correct reasoning over visual premises, requiring strict logical consistency unlike traditional metrics that allow partial correctness. It requires that every premise in a reasoning sequence be validated as correct for the entire sequence to be considered accurate, reflecting the principles of formal logic where any single error can invalidate an entire proof. This follows from the nature of deductive reasoning, where conclusions are only as valid as their weakest premise. The Depth of Correct Premises ($DCP$) metric complements $Acc_{VPP}$ by providing insight into the extent of errors within incorrect sequences.
3.3.1 Accuracy of Proven Premises
Let $S$ denote the total number of sequences, $P_{s}$ the number of premises in sequence $s$, and $C_{s,p}$ a binary indicator of correctness for premise $p$ in sequence $s$. The $Acc_{VPP}$ metric is defined as:
$Acc_{VPP}=\frac{1}{S}\sum_{s=1}^{S}\left(\prod_{p=1}^{P_{s}}C_{s,p}\right)$
This metric assigns a maximum value of 1 for sequences where all premises are true and 0 if any premise is incorrect. The multiplicative approach reflects the logical "AND" operation in formal proofs, where the truth value of the conclusion depends on all components being true. If a model predicts a final conclusion as True, all intermediate premises in its reasoning sequence must also be True; otherwise, an incorrect intermediate premise indicates that the model relied on an incorrect bias to reach the correct conclusion.
3.3.2 Depth of Correct Premises
To assess how well a model reasons through incomplete sequences, we use the Depth of Correct Premises ($DCP$) metric:
${DCP}=\frac{1}{S-S_{correct}}\sum_{s=1}^{S}\left(\frac{\sum_{p=1}^{P_{s}}C_{s,p}}{P_{s}}\right),\text{where }\prod_{p=1}^{P_{s}}C_{s,p}=0$
Here, $S_{correct}$ represents the number of sequences where all premises are correct. The $DCP$ metric first identifies incorrect sequences and normalizes the sum of correct premises by the total number of premises in each chain. This provides a measure of how far a model correctly reasons within sequences that are not fully accurate, highlighting areas for potential improvement in model training and reasoning formulation.
While metrics like NDCG and MRR offer measures for structured output, they do not assess the depth of correctness within reasoning sequences. The $DCP$ metric captures how much of the reasoning chain is correct even when the overall sequence is not, providing a nuanced understanding of the model’s capacity to handle extended reasoning tasks.
4 Experimentation
We conduct all our experiments using a 12 layer encoder-decoder transformer model with a resolution preserving image tokenizer as introduced in Pix2Struct. We employ the variant pretrained with math and chart question answering following Matcha. All our experiments are conducted using 4XNvidia A6000 (24G) machines with a batch size of 2. The model uses a fixed context length of 2048 tokens. And inference with a single text-image pair costs about $\sim$5G memory. ChartPrem model is trained using only structural premises, previous question answer strings and groundtruth annotation-json from the original UB-PMC task. 1 full epoch with chart qa (1.5mil) + challenge json (30k) + structure premises(52k) takes around 2 weeks in our setting. We train for 150k iterations(see supplementary) and use all types of the premises for evaluation (158k) of our proposed task, using a subset of the extensively generated data premises(60k). We hope by publicly releasing the full dataset, models and evaluation code, it will stimulate the community for further exploration especially using the extensive premises for instruction fine-tuning LVLM’s and lead to improvement of models to parse large scale scientific publication repositories.
4.1 Dense Chart Parsing
We compare our model’s performance on dense chart parsing tasks by using structured premises for structure prediction, such as querying ’What is the type of chart?’ or ’What is the title of the dependent axis?’ instead of conventional image classification. We compare with the results of the corresponding tasks from the previous challenge[11], and in the column ‘Direct Prediction’ we report results as reported in the challenge report by task specific vision models. The next column is the Matcha-base model in a zero shot setting, and the next column ‘Matcha-FT’ is the same model when trained only on the RealCQA QA pairs. The next column, ‘ChartPrem’ is our proposed model trained further on the SP’s we created. We provide exhaustive list of our queries in the appendix. These queries are complementary to the underlying binary SP’s that we created for the VPP task and have text answers. Thus, training on SP’s improves performance on previous dense chart parsing tasks as compared to direct pixel level predictions. While zero-shot Matcha is only able to generalize sufficiently to three chart properties categorical labels, logarithmic axis and presence of legend, on fine-tuning the performance improves but the vision based task specific models still outperform. Only on further training with the SP’s we see considerable improvement. The worst performance is for dependent axis title which can include multiple complex math symbols used in scientific charts which the model might not have seen as much due to the text heavy pre-training datasets. The second worst performance is for categorical x-tick labels and this is due to the complex grouped and stacked bar charts which might again have math-symbol intensive labels and at times are quite cluttered having 40-50 tilted labels on a single chart.
Chart Component Task Direct Prediction Matcha Matcha(FT) ChartPrem(Ours) Evaluation Metric
Chart Type 94.63 35.21 80.96 98.72 F1 - Precision/recall
Dependent(Y) Axis Title 75.62 21.85 47.82 77.80 Text Matching
Y-Min Value 73.42 34.61 70.1 97.59 Absolute Value
Y-Max Value 62.22 24.74 54.63 96.5 Absolute Value
Independent(X) Axis Title 85.62 42.3 79.34 92.38 Text Matching
X-Min Value 73.42 24.61 62.77 96.94 Absolute Value
X-Max Value 62.22 34.74 57.02 95.31 Absolute Value
Categorical X-Tick Labels 68.83 33.78 58.76 79.74 1:1 Accuracy
Is Categorical or Not - 82.97 91.48 100 Binary
Is Logarithmic or Not - 84.65 93.24 100 Binary
Is Legend Present or Not - 77.32 90.66 100 Binary
Number of Data Series - 38.43 56.57 82.61 Absolute Value
Legend Name 82.93 47.92 73.86 91.27 1:1 Accuracy
Table 2: Dense Chart Parsing on UB-PMC (Underlined models are Zero-Shot, rest Fine-Tuned.)
4.2 NLP QA
"We evaluate our premises-trained model against the standard benchmark for question answering tasks from RealCQA, using the RQA9357 subset.We evaluate the task of chart question answering through multiple views. Figure 4 shows the delta performance of each model with comparison to zero-shot matcha-base. These results are presented in Table 3, where the first column describes the category of QA, the rest are performances by different models. These are all reported as accuracy percent over total QA pairs in that category except ranked lists which are evaluated using nDCG@10.
Evaluation Set (# Total-QA) VL-T5 CRCT UniChart Matcha Matcha(FT) ChartPrem(Ours)
Total Accuracy % 367,139 31.06 18.80 26.75 25.97 32.10 44.62
Answer Type
String Answer % 19,525 30.68 3.23 0.88 2.47 29.50 83.97
Numerical Answer % 115,391 14.87 31.58 0.83 4.01 13.39 15.68
Ranked Answer nDCG10 16,389 0.0246 0.0286 0.0113 0.0088 0.270 0.322
UnRanked Answer % 44,702 0.48 1.24 0.14 0.20 16.03 28.11
Binary Answer % 171,132 52.75 18.07 51.53 52.54 56.19 67.95
Question Type
Structural Question % 48,306 43.52 14.98 21.40 19.85 42.41 83.89
Retrieval Question % 8,220 58.77 31.31 24.72 14.20 50.82 62.44
Reasoning Question % 310,613 29.37 19.60 27.64 27.71 30.89 38.84
Chart Type
Line Chart % 115,899 38.24 19.06 33.51 32.67 39.78 50.72
Vertical Bar Chart % 178,740 28.79 15.06 22.99 22.95 29.60 39.69
Horizontal Bar Chart % 46,214 25.42 29.17 20.58 17.19 25.56 35.45
Scatter Chart % 4,371 28.29 8.07 16.09 18.19 36.09 81.81
Vertical Box Chart % 21,915 24.06 11.84 36.93 41.99 52.86 64.52
Table 3: NLP-QA on RealCQA (Underlined models are Zero-Shot, rest Fine-Tuned.)
Premise Category Accuracy FOLS $ACC_{VPP}$ DCP Total Train Total Test(Used)
Overall 0.46 0.2 0.395 2,156,628 548,564
Structural Premises 0.88 0.34 0.73 52,665 14,778(all)
Data Premises 0.56 0.21 0.43 1,609,374 405,541(60k)
Reasoning Premises 0.27 0.14 0.38 328,293 78,999(all)
Math Premises 0.13 0.07 0.04 8,301 4,912(all)
Table 4: FOLS and VPP for Real CQA, reported counts are for True premises, with False we have 4x
Figure 4: Taking Zero Shot Matcha as baseline comparison over NLP-QA (Best viewed digital zoom and color)
The VL-T5 and CRCT results are of fine-tuned models as provided by the authors of RealCQA. The UniChart and Matcha models are evaluated in zero shot setting. The Matcha(FT) and ChartPrem are as previously described the baseline trained only on question answers and the proposed training with premises.
The first set is based on the answer type. While we see VL-T5 and Matcha models perform well on binary answers the CRCT model which handles numeric and string answers through regression and classification separately performs better on these. All models perform best on binary answers and worst on ranked lists. Fine tuning using only question answer pairs improves results much more for all types and only slightly for the binary answer type. This seems to suggest that while binary answers are more easily generalizable, the model needs to visually adapt to the scientific domain to perform better on the rest and that while the zero shot model has enough capacity to represent language to answer binary type questions, it still has to adapt more to the complex notations involved in scientific charts to correctly parse the math and text in these images. The ChartPrem version trained with the premises shows the most gains in string type answers and also unranked lists, suggesting the considerable importance of the chart structure and step wise reasoning.
The next categorisation is by question type. It is interesting to note the zero-shot Matcha model has better performance than CRCT on structural and reasoning questions owing to their larger scale pre-training dataset, but lags in retrieval due to the visual domain shift of scientific charts. Both zero shot models Unichart and Matcha perform better on reasoning type questions than structural and retrieval based questions. This suggests these models are capable of leveraging other inductive biases present in the data and perform well without actually performing ‘reasoning’ as depicted by only looking at this metric.
Model Accuracy $\Delta$ Correct QA
CP 44.62 0
CP + SD 44.77 551$\uparrow$
CP + CD 44.84 808$\uparrow$
CP + SD + CD 44.96 1249$\uparrow$
Table 5: Ablation NLP-QA on RealCQA
The last categorization is based on the chart type. Owing to the natural PMC distribution the RQA dataset has more QA on line($\sim 31.56\%$) and vertical bar ($\sim 48.68\%$) than horizontal bar ($\sim 12.58\%$), scatter ($\sim 1.19\%$) and box($\sim 5.96\%$) charts. The zero-shot perform best on line and vertical bar (seen), also vertical box (new) and worst on horizontal bar, scatter (new). QA training improves under-represented chart types accuracy. Chart premises training shows the most gains with scatter type plots. he challenge with scatter plots[16] lies in distinguishing between axis ticks, legends, and point markers, which are vissually same but semantically different depending on location. By learning the global context of chart structure, ChartPrem better differentiates between local data points and the overall chart structure.
We conduct an ablation study by employing early fusion of trained vision encoders for chart structure (CACHED[16]) and chart data extraction (SpaDen [17]). We extract features from the penultimate layer, then use a simple MLP to transform, normalize, and add them as additional tokens for the ChartPrem input transformer encoder. Table 5 shows only nominal gains, suggesting that while these models do have specialized features, they provide only limited useful signals. Further research is needed to explore more effective fusion techniques or alternative representations like GNN math embeddings[25]. Currently, the marginal gains do not justify the effort required to train three individual models.
4.3 Visual Premise Proving
We evaluate the ChartPrem model on our VPP metrics.
4.3.1 FOL Solver
Our first analysis is based on the models capability to prove individual FOL premises. As shown in the first column of Table 4. Overall, the ChartPrem model is able to solve 46% of binary FOL performing best on structure and worst on Math. this is expected since the current model has not be trained on other premise types. Surprisingly it still get 56% of data premises correct suggesting this style of training/inference can be effective in extracting tabular data used to plot the chart.
4.3.2 Chain of Reasoning Evaluation
In Table 4 we observed an overall $Acc_{VPP}$ of 0.2 and a $DCP$ of 0.395 for our model which has an impressive 68% accuracy on binary reasoning QA as discussed previously.
The $Acc_{VPP}$ value of 0.2 indicates that only 20% of the sequences in the dataset were completely correct, where every premise within these sequences was validated accurately. This relatively low score suggests that the model struggles to completely validate all premises in most sequences. Even for the high 88% FOLS accuracy of SP’s the sequential correctness is only 34%. On the other hand, the $DCP$ value of 0.395 reveals that, on average, approximately 39.5% of the premises are correctly validated in sequences where not all premises are correct. This metric is particularly revealing, as it quantifies the extent to which the model can successfully navigate through part of the reasoning chain before making an error. A DCP less than 0.5 indicates that, in sequences with errors, the model tends to fail before reaching the halfway point of the premise sequence on average. This means a specific weaknesses to handle more extended chains of reasoning.
Together, these metrics suggest that while the model has some capability to process and validate individual premises, it struggles with consistency and completeness in more extended reasoning tasks.
5 Conclusion
We curated over 10 million text descriptors for a real-world chart dataset and demonstrated that high reasoning QA accuracy alone does not fully measure a model’s reasoning capabilities. Certain limitations also remain. First, RealCQA is the only source providing real-world charts with the necessary annotations. Models trained on our premises could be used to generate logic sequences for other chart datasets, which would still require logic verification. Second, our focus on Charts and the constraints set by first-order logic, while necessary due to the nascent stage of visual reasoning research, limits the direct applicability of our approach to broader vision tasks. Once the domain has sufficient traction over chart/table/document QA one could envision identification of visual logic sequences in the wild. Lastly, our work emphasizes the challenge in visual reasoning due to visual complexity but does not address the full spectrum of natural language variation. We do paraphrase premises but the original questions still remain template based and addressing both visual and language complexities would exceed the scope of a single study.